OpenStudy (anonymous):

What is the slope of a line perpendicular to the line with equation y = 6x + 2?

OpenStudy (abb0t):

Perpendicular is basically the reciprocal of the slope. When you have: \(\sf \color{Red}{y=mx+b}\) m = slope

OpenStudy (anonymous):

m=-1/6

OpenStudy (anonymous):

slope = 6

OpenStudy (abb0t):

slope is: \(\sf \frac{6}{1}\)

OpenStudy (anonymous):

\[A. - \frac{ 1 }{ 6 } B. 2 C. -6 . 6 \]

OpenStudy (dan815):